# A Survey of Vibe Coding with Large Language Models

### v1

@@ 7.2.1 Unit Test Execution Results
Unit test execution results serve as critical feedback mechanisms allowing coding agents to verify generated code meets requirements [528, 181, 175, 509]. Results provide varying detail levels: simple correctness feedback indicating pass/fail through binary signals, and detailed execution results including runtime errors and specific test failure information. Test-Driven Development (TDD) has proven its merit, requiring developers to write tests before the functional code [529, 253, 530]. Test validation frameworks like PyTest, unittest, and JUnit serve as primary mechanisms generating feedback revealing runtime errors [524, 531, 243].
Unit test results serve multiple applications: reinforcement learning approaches leveraging execution feedback, search-based debugging where algorithms like BESTER use execution feedback with self-reflection for automated repairs [275], code sample filtering using test outcomes with CodeT achieving 65.8% pass@1 on HumanEval, performance optimization analyzing execution time data [176, 170], multi-agent systems employing execution feedback [171, 532, 13], and test-time scaling strategies generating multiple candidates.
Integration methodologies include reinforcement learning approaches fine-tuning models using test execution outcomes assigning rewards based on results using algorithms like REINFORCE or PPO optimizing model behavior through iterative feedback loops [245, 246], code selection and filtering using execution feedback to identify best solutions with AlphaCode generating millions of programs filtering incorrect solutions based on test outcomes [269], sophisticated approaches like CodeT automatically generating test cases performing dual execution agreement analysis considering both output consistency and cross-sample agreement [176], and structured validation frameworks implementing unit testing as iterative refinement mechanisms where agents continuously generate test cases, evaluate outputs against functional requirements, and refine code until all tests pass.
Despite promising applications, significant challenges limit effectiveness: unreliability of AI-generated unit tests where code refinement depends entirely on generated unit test correctness, with studies revealing models generate unreliable self-tests in abundance potentially misleading self-improvement processes and exacerbating program errors creating circular dependency problems [172, 173], scalability constraints as generating and executing multiple test cases for each code candidate requires significant computational resources particularly when systems generate thousands of code samples, and limited test coverage as automatically generated tests may not capture all edge cases or complex scenarios potentially allowing buggy code to pass initial validation while failing in real-world deployment scenarios.
@@ 7.2.2 Integration Test Feedback
Integration test feedback involves executing automated tests against LLM-generated code and using resulting feedback to guide subsequent iterations, building upon traditional software testing but adapting for unique challenges of AI-generated code [492]. The fundamental premise relies on creating closed-loop systems where coding agents learn from execution feedback rather than relying solely on static analysis or human review.
Development of integration testing frameworks has evolved to address unique challenges of testing LLM-powered systems [26], with OpenHands framework representing notable advancement pioneering end-to-end agent test frameworks combining traditional integration testing with foundation model mocking techniques [183, 179, 533].
Practical implementation relies on sophisticated multi-agent architectures where test executor agents serve as central coordinators managing interaction between code generation and test validation activities. Effectiveness varies significantly based on type and structure of feedback, with research demonstrating structured feedback formats achieving superior results, with test feedback leading to highest repair success rates [492]. The iterative nature plays crucial roles with multiple rounds enhancing repair performance, though research indicates diminishing returns after two or three iterations [492].
Filtering and selection of generated code samples involves several established approaches. Key methods include test-based filtering using unit test outcomes [269], dual execution agreement evaluating code based on output consistency [176], verification-based reranking training specialized verifiers [534, 535], semantic execution analysis using execution results to capture semantic features [167], and probabilistic marginalization combining programs with identical execution results [176].
@@ 7.2.3 Runtime Error and Exception Handling
Research frameworks have systematically categorized runtime errors when LLMs generate code identifying five primary error types: syntax errors, parameter and attribute errors, output type errors, logical answer errors, and timeout and runtime errors [516]. Beyond execution-level errors, comprehensive taxonomies identify 36 exception types across 12 agent artifacts distinguishing failures across reasoning, planning, and execution phases [74, 74].
Research has developed several approaches for detecting and analyzing runtime errors: simplest approaches using binary feedback providing only pass/fail indicators [428], sophisticated systems offering failed tests feedback including expected versus actual values with runtime error tracebacks [429], advanced frameworks providing failed and passed tests feedback with comprehensive input/output information [167], sophisticated debugging approaches using control flow analysis and variable tracking with systems like LDB [429], and emerging print debugging and reflection-based methods [111, 428].
Several specialized frameworks provide comprehensive exception management: knowledge-driven approaches where Knowledge-driven Prompt Chaining decomposes code generation into AI chains with iterative check-rewrite steps [178], multi-agent frameworks like Seeker using specialized agents [180], PyCapsule employing programmer agents and debugging alongside executor agents enhanced by error handlers [174], runtime error recovery systems like Healer with GPT-4 successfully handling 88.1% of AttributeErrors [182], and iterative self-correction mechanisms using structured feedback loops [177].
Self-reflection represents critical capability for LLM-based coding agents defined as processes where models review earlier outputs and attempt to identify logical flaws. When given buggy programs with execution traces, errors, or test results as input, LLMs produce debugging instructions through self-reflection [118]. Dynamic self-correction through execution feedback represents sophisticated approaches where systems append diagnostic information to original prompts [177]. Modern frameworks have developed increasingly sophisticated debugging methodologies including print debugging methods, reflection-based approaches, hierarchical debugging tools, and advanced systems like CYCLE [111, 163].
@@ 7.3.1 Interactive Requirement Clarification
Large Language Models have demonstrated remarkable capabilities in automatically generating code from natural language requirements [35], though critical gaps exist between current LLM behavior and human problem-solving approaches when dealing with ambiguous requirements [188, 536, 497]. While human developers typically seek additional information through interactive clarification, current LLM-based code generation approaches lack mechanisms for clarifying unclear requirements, with empirical studies revealing state-of-the-art Code LLMs generate code outputs in over 63% of ambiguous scenarios without seeking necessary clarifications [188]. This limitation has significant implications compounded by the fact that 72% of software defects in production environments originate from misunderstood requirements [536, 184, 471].
Detection of ambiguous requirements has emerged as critical first steps with the most widely adopted approach being code consistency checking pioneered by ClarifyGPT which determines whether given requirements are ambiguous by generating multiple code solutions [188]. Despite methodological advances, empirical findings reveal significant limitations with systematic evaluations showing state-of-the-art models struggle to distinguish between well-specified and underspecified instructions [243, 112].
Generation of effective clarification questions represents core technical challenges with foundational approaches established by ClarifyGPT using prompt-based question generation, demonstrating significant improvements elevating GPT-4’s performance from 70.96% to 80.80% on MBPP-sanitized benchmarks [188]. To address limitations, fine-tuning approaches have emerged with ClarifyCoder representing notable advances using synthetic data generation and instruction-tuning [185, 536, 188, 113].
Evaluation of interactive requirement clarification systems has evolved toward systematic assessment frameworks. Empirical findings reveal significant performance gaps—while models struggle to distinguish between well-specified and underspecified instructions, when they do interact for underspecified inputs they effectively obtain vital information from users leading to significant improvements with evaluations showing average absolute improvement of 45.97% in pass@1 code generation accuracy within 5 user interactions [243, 112].
@@ 7.3.2 Code Review Feedback
Development of large language models has created urgent needs to align these powerful systems with human intentions and preferences, with the challenge lying in effectively utilizing qualitative human feedback [223]. Foundations of this field can be traced to preference-based reinforcement learning (PbRL), with major breakthroughs coming with Christiano et al.’s introduction of RLHF [537].
@@
RLHF framework follows structured approaches beginning with SFT on high-quality demonstration data, with core innovation lying in training reward models to approximate human preferences from preference data where humans indicate preferred responses between pairs of options serving as foundations for subsequent reinforcement learning optimization [189]. The field gained significant momentum when scaled to large language models [538] with InstructGPT demonstrating potential of applying RLHF to large-scale models like GPT-3 enabling alignment with user instructions across wide ranges of tasks [8], while DPO has emerged as notable simplification allowing language models to be trained directly on preference data without explicit reward modeling or reinforcement learning optimization stages [2, 70, 186, 68, 227].
@@
Core RLHF methodology consists of three distinct stages: processes begin with supervised fine-tuning on high-quality demonstration data, second stages focus on reward modeling training reward models to approximate human preferences based on Bradley-Terry models, and final stages involve reinforcement learning optimization where models are refined using RL algorithms like PPO [539, 2, 190, 68]. Despite effectiveness, traditional RLHF pipelines are notably complex and computationally expensive [68].
@@
Traditional RLHF pipelines suffer from inherent complexity and instability creating practical barriers [68]. Major practical constraints include substantial costs and efforts required for human annotation often requiring thousands of human comparisons to train reliable reward models [191, 540]. Scalability challenges become even more pronounced with complex tasks as dialogues become longer spanning multiple turns greatly increasing annotation difficulty [191].
@@
Given these limitations, complexity and instability of traditional RLHF has motivated development of more direct approaches with most significant breakthrough coming with DPO eliminating explicit reward modeling stages and directly optimizing policies using preference data leveraging mathematical relationships between optimal policies and reward functions under Bradley-Terry preference models enabling preference optimization through simple classification objectives that are stable, performant, and computationally lightweight [68]. Beyond DPO, researchers have explored alternative frameworks. Preference Ranking Optimization (PRO) addresses limitations of RLHF by extending pairwise preference comparisons to accommodate human preference rankings of any length directly fine-tuning language models through iterative contrast learning [541]. Nash Learning from Human Feedback (NLHF) offers novel approaches learning preference models conditioned on two inputs pursuing strategies consistently generating responses more preferred than competing policies [542]. Diffusion-DPO demonstrates effectiveness in text-to-image generation [543, 544].
@@
Substantial costs required for human annotation in traditional RLHF have motivated researchers to explore AI-generated feedback as scalable alternatives with RLAIF introducing frameworks where AI-generated feedback replaces human annotations with off-the-shelf LLMs serving as feedback models to evaluate responses based on predefined principles demonstrating performance comparable to traditional RLHF methods across tasks like summarization and dialogue generation [190, 2, 538, 187]. AutoPM represents another approach employing LLMs to automatically generate preference data by eliciting pairwise comparisons based on helpfulness, honesty, and harmlessness criteria enabling training of preference models without extensive human supervision [190, 539].
@@ 7.4.1 Self-Evaluation and Critique
Foundational architecture of self-refinement in LLMs centers around iterative feedback loops mirroring human revision processes with most notable framework Self-Refine enabling LLMs to improve outputs through structured three-step processes: generating initial responses, prompting models to critique own outputs, and refining responses based on feedback [210, 105, 545]. Processes follow key steps: initial generation, self-critique, refinement, and iteration, with notably Self-Refine requiring no additional training data or fine-tuning [105, 72]. Multiple variants have emerged including Recursively Criticizes and Improves (RCI), CRITIC, and Self-Correct [192, 210].
Executable nature of code provides unique advantages for self-refinement in programming tasks as execution results offer immediate and objective feedback [286, 105, 201]: Self-Debugging equipping models with ability to autonomously detect and repair errors, Self-Edit incorporating dedicated fault detection phases, CodeChain introducing self-revision mechanisms, Self-Collaboration using simulated internal dialogue, LeTI redefining code generation as interactive dialogue-driven processes, and OpenCodeInterpreter unifying generation, execution, and refinement.
Self-evaluation represents critical components encompassing models’ ability to assess quality of own generated outputs including reasoning steps and confidence scores with multiple techniques having emerged including self-consistency, self-correction, self-evolution, and self-feedback [193]. Self-critique and feedback generation processes involve models acting as critics analyzing own outputs [105, 114]. However, critical analysis reveals important limitations with research indicating self-correction works well primarily in tasks that can use reliable external feedback, and large-scale fine-tuning appears necessary to enable effective self-correction [193, 546].
Critique generation processes involve models analyzing own outputs and providing structured suggestions for improvement [105]. Advanced critique frameworks implement multi-dimensional evaluation systems with carefully designed prompts instructing models to assess outputs based on specific criteria [114, 105, 201]. Performance gains from self-refinement methods have been substantial including GPT-4 achieving gains of 8.7 points in code optimization and 13.9 points in code readability when using Self-Refine prompting [210].

### v2

@@ 7.2.1 Unit Test Execution Results
Unit test execution results serve as critical feedback mechanisms allowing coding agents to verify generated code meets requirements [yang2024evaluation, alshahwan2024automated, pan2025aster, dolcetti2024helping]. Results provide varying detail levels: simple correctness feedback indicating pass/fail through binary signals, and detailed execution results including runtime errors and specific test failure information. Test-Driven Development (TDD) has proven its merit, requiring developers to write tests before the functional code [beck2003test, mathews2024test, cui2025tests]. Test validation frameworks like PyTest, unittest, and JUnit serve as primary mechanisms generating feedback revealing runtime errors [lai2022natural, gu2024testart, dong2025survey].
Unit test results serve multiple applications: reinforcement learning approaches leveraging execution feedback, search-based debugging where algorithms like BESTER use execution feedback with self-reflection for automated repairs [goues2019automated], code sample filtering using test outcomes with CodeT achieving 65.8% pass@1 on HumanEval, performance optimization analyzing execution time data [chen2022codet, peng2024perfcodegen], multi-agent systems employing execution feedback [bouzenia2024you, s2025multi, zhang2024autocoderover], and test-time scaling strategies generating multiple candidates.
Integration methodologies include reinforcement learning approaches fine-tuning models using test execution outcomes assigning rewards based on results using algorithms like REINFORCE or PPO optimizing model behavior through iterative feedback loops [le2022coderl, shojaee2023execution], code selection and filtering using execution feedback to identify best solutions with AlphaCode generating millions of programs filtering incorrect solutions based on test outcomes [li2022competition], sophisticated approaches like CodeT automatically generating test cases performing dual execution agreement analysis considering both output consistency and cross-sample agreement [chen2022codet], and structured validation frameworks implementing unit testing as iterative refinement mechanisms where agents continuously generate test cases, evaluate outputs against functional requirements, and refine code until all tests pass.
Despite promising applications, significant challenges limit effectiveness: unreliability of AI-generated unit tests where code refinement depends entirely on generated unit test correctness, with studies revealing models generate unreliable self-tests in abundance potentially misleading self-improvement processes and exacerbating program errors creating circular dependency problems [liu2025llm, wang2025projecttest], scalability constraints as generating and executing multiple test cases for each code candidate requires significant computational resources particularly when systems generate thousands of code samples, and limited test coverage as automatically generated tests may not capture all edge cases or complex scenarios potentially allowing buggy code to pass initial validation while failing in real-world deployment scenarios.
@@ 7.2.2 Integration Test Feedback
Integration test feedback involves executing automated tests against LLM-generated code and using resulting feedback to guide subsequent iterations, building upon traditional software testing but adapting for unique challenges of AI-generated code [dai2025feedbackeval]. The fundamental premise relies on creating closed-loop systems where coding agents learn from execution feedback rather than relying solely on static analysis or human review.
Development of integration testing frameworks has evolved to address unique challenges of testing LLM-powered systems [yang2023intercode], with OpenHands framework representing notable advancement pioneering end-to-end agent test frameworks combining traditional integration testing with foundation model mocking techniques [schultz2025potential, zhang2024llm, nong2024automated].
Practical implementation relies on sophisticated multi-agent architectures where test executor agents serve as central coordinators managing interaction between code generation and test validation activities. Effectiveness varies significantly based on type and structure of feedback, with research demonstrating structured feedback formats achieving superior results, with test feedback leading to highest repair success rates [dai2025feedbackeval]. The iterative nature plays crucial roles with multiple rounds enhancing repair performance, though research indicates diminishing returns after two or three iterations [dai2025feedbackeval].
Filtering and selection of generated code samples involves several established approaches. Key methods include test-based filtering using unit test outcomes [li2022competition], dual execution agreement evaluating code based on output consistency [chen2022codet], verification-based reranking training specialized verifiers [zhang2024generative, zhang2023coder], semantic execution analysis using execution results to capture semantic features [ni2025viscoder], and probabilistic marginalization combining programs with identical execution results [chen2022codet].
@@ 7.2.3 Runtime Error and Exception Handling
Research frameworks have systematically categorized runtime errors when LLMs generate code identifying five primary error types: syntax errors, parameter and attribute errors, output type errors, logical answer errors, and timeout and runtime errors [li2024redo]. Beyond execution-level errors, comprehensive taxonomies identify 36 exception types across 12 agent artifacts distinguishing failures across reasoning, planning, and execution phases [yao2022react, yao2022react].
Research has developed several approaches for detecting and analyzing runtime errors: simplest approaches using binary feedback providing only pass/fail indicators [chen2025revisit], sophisticated systems offering failed tests feedback including expected versus actual values with runtime error tracebacks [zhong2024ldb], advanced frameworks providing failed and passed tests feedback with comprehensive input/output information [ni2025viscoder], sophisticated debugging approaches using control flow analysis and variable tracking with systems like LDB [zhong2024ldb], and emerging print debugging and reflection-based methods [ashrafi2025enhancing, chen2025revisit].
Several specialized frameworks provide comprehensive exception management: knowledge-driven approaches where Knowledge-driven Prompt Chaining decomposes code generation into AI chains with iterative check-rewrite steps [ren2023from], multi-agent frameworks like Seeker using specialized agents [zhang2024seeker], PyCapsule employing programmer agents and debugging alongside executor agents enhanced by error handlers [adnan2025large], runtime error recovery systems like Healer with GPT-4 successfully handling 88.1% of AttributeErrors [sun2024llm], and iterative self-correction mechanisms using structured feedback loops [rath2025structured].
Self-reflection represents critical capability for LLM-based coding agents defined as processes where models review earlier outputs and attempt to identify logical flaws. When given buggy programs with execution traces, errors, or test results as input, LLMs produce debugging instructions through self-reflection [jin2024rgd]. Dynamic self-correction through execution feedback represents sophisticated approaches where systems append diagnostic information to original prompts [rath2025structured]. Modern frameworks have developed increasingly sophisticated debugging methodologies including print debugging methods, reflection-based approaches, hierarchical debugging tools, and advanced systems like CYCLE [ashrafi2025enhancing, ding2024cycle].
@@ 7.3.1 Interactive Requirement Clarification
Large Language Models have demonstrated remarkable capabilities in automatically generating code from natural language requirements [chen2021evaluating], though critical gaps exist between current LLM behavior and human problem-solving approaches when dealing with ambiguous requirements [mu2024clarifygpt, darji2025curiosity, fang2024large]. While human developers typically seek additional information through interactive clarification, current LLM-based code generation approaches lack mechanisms for clarifying unclear requirements, with empirical studies revealing state-of-the-art Code LLMs generate code outputs in over 63% of ambiguous scenarios without seeking necessary clarifications [mu2024clarifygpt]. This limitation has significant implications compounded by the fact that 72% of software defects in production environments originate from misunderstood requirements [darji2025curiosity, barke2023grounded, ross2023programmer].
Detection of ambiguous requirements has emerged as critical first steps with the most widely adopted approach being code consistency checking pioneered by ClarifyGPT which determines whether given requirements are ambiguous by generating multiple code solutions [mu2024clarifygpt]. Despite methodological advances, empirical findings reveal significant limitations with systematic evaluations showing state-of-the-art models struggle to distinguish between well-specified and underspecified instructions [dong2025survey, fakhoury2024llm].
Generation of effective clarification questions represents core technical challenges with foundational approaches established by ClarifyGPT using prompt-based question generation, demonstrating significant improvements elevating GPT-4’s performance from 70.96% to 80.80% on MBPP-sanitized benchmarks [mu2024clarifygpt]. To address limitations, fine-tuning approaches have emerged with ClarifyCoder representing notable advances using synthetic data generation and instruction-tuning [marozzo2025iterative, darji2025curiosity, mu2024clarifygpt, lahiri2022interactive].
Evaluation of interactive requirement clarification systems has evolved toward systematic assessment frameworks. Empirical findings reveal significant performance gaps—while models struggle to distinguish between well-specified and underspecified instructions, when they do interact for underspecified inputs they effectively obtain vital information from users leading to significant improvements with evaluations showing average absolute improvement of 45.97% in pass@1 code generation accuracy within 5 user interactions [dong2025survey, fakhoury2024llm].
@@ 7.3.2 Code Review Feedback
Development of large language models has created urgent needs to align these powerful systems with human intentions and preferences, with the challenge lying in effectively utilizing qualitative human feedback [kaufmann2024survey]. Foundations of this field can be traced to preference-based reinforcement learning (PbRL), with major breakthroughs coming with Christiano et al.’s introduction of RLHF [christiano2017deep].
@@
RLHF framework follows structured approaches beginning with SFT on high-quality demonstration data, with core innovation lying in training reward models to approximate human preferences from preference data where humans indicate preferred responses between pairs of options serving as foundations for subsequent reinforcement learning optimization [schulman2017proximal]. The field gained significant momentum when scaled to large language models [lee2023rlaif] with InstructGPT demonstrating potential of applying RLHF to large-scale models like GPT-3 enabling alignment with user instructions across wide ranges of tasks [achiam2023gpt], while DPO has emerged as notable simplification allowing language models to be trained directly on preference data without explicit reward modeling or reinforcement learning optimization stages [ouyang2022training, stiennon2020learning, zaheer2017deep, rafailov2023direct, xiao2024comprehensive].
@@
Core RLHF methodology consists of three distinct stages: processes begin with supervised fine-tuning on high-quality demonstration data, second stages focus on reward modeling training reward models to approximate human preferences based on Bradley-Terry models, and final stages involve reinforcement learning optimization where models are refined using RL algorithms like PPO [yin2025from, ouyang2022training, niekerk2025post, rafailov2023direct]. Despite effectiveness, traditional RLHF pipelines are notably complex and computationally expensive [rafailov2023direct].
@@
Traditional RLHF pipelines suffer from inherent complexity and instability creating practical barriers [rafailov2023direct]. Major practical constraints include substantial costs and efforts required for human annotation often requiring thousands of human comparisons to train reliable reward models [kiruluta2025history, kopf2023openassistant]. Scalability challenges become even more pronounced with complex tasks as dialogues become longer spanning multiple turns greatly increasing annotation difficulty [kiruluta2025history].
@@
Given these limitations, complexity and instability of traditional RLHF has motivated development of more direct approaches with most significant breakthrough coming with DPO eliminating explicit reward modeling stages and directly optimizing policies using preference data leveraging mathematical relationships between optimal policies and reward functions under Bradley-Terry preference models enabling preference optimization through simple classification objectives that are stable, performant, and computationally lightweight [rafailov2023direct]. Beyond DPO, researchers have explored alternative frameworks. Preference Ranking Optimization (PRO) addresses limitations of RLHF by extending pairwise preference comparisons to accommodate human preference rankings of any length directly fine-tuning language models through iterative contrast learning [song2024preference]. Nash Learning from Human Feedback (NLHF) offers novel approaches learning preference models conditioned on two inputs pursuing strategies consistently generating responses more preferred than competing policies [munos2023nash]. Diffusion-DPO demonstrates effectiveness in text-to-image generation [weng2024navigating, wallace2023diffusion].
@@
Substantial costs required for human annotation in traditional RLHF have motivated researchers to explore AI-generated feedback as scalable alternatives with RLAIF introducing frameworks where AI-generated feedback replaces human annotations with off-the-shelf LLMs serving as feedback models to evaluate responses based on predefined principles demonstrating performance comparable to traditional RLHF methods across tasks like summarization and dialogue generation [niekerk2025post, ouyang2022training, lee2023rlaif, chai2024ma]. AutoPM represents another approach employing LLMs to automatically generate preference data by eliciting pairwise comparisons based on helpfulness, honesty, and harmlessness criteria enabling training of preference models without extensive human supervision [niekerk2025post, yin2025from].
@@ 7.4.1 Self-Evaluation and Critique
Foundational architecture of self-refinement in LLMs centers around iterative feedback loops mirroring human revision processes with most notable framework Self-Refine enabling LLMs to improve outputs through structured three-step processes: generating initial responses, prompting models to critique own outputs, and refining responses based on feedback [sahoo2024systematic, madaan2023self, he2025can]. Processes follow key steps: initial generation, self-critique, refinement, and iteration, with notably Self-Refine requiring no additional training data or fine-tuning [madaan2023self, wei2022chain]. Multiple variants have emerged including Recursively Criticizes and Improves (RCI), CRITIC, and Self-Correct [gou2023critic, sahoo2024systematic].
Executable nature of code provides unique advantages for self-refinement in programming tasks as execution results offer immediate and objective feedback [zheng2024opencodeinterpreter, madaan2023self, shinn2023reflexion]: Self-Debugging equipping models with ability to autonomously detect and repair errors, Self-Edit incorporating dedicated fault detection phases, CodeChain introducing self-revision mechanisms, Self-Collaboration using simulated internal dialogue, LeTI redefining code generation as interactive dialogue-driven processes, and OpenCodeInterpreter unifying generation, execution, and refinement.
Self-evaluation represents critical components encompassing models’ ability to assess quality of own generated outputs including reasoning steps and confidence scores with multiple techniques having emerged including self-consistency, self-correction, self-evolution, and self-feedback [mehandru2025bioagents]. Self-critique and feedback generation processes involve models acting as critics analyzing own outputs [madaan2023self, ho2025self]. However, critical analysis reveals important limitations with research indicating self-correction works well primarily in tasks that can use reliable external feedback, and large-scale fine-tuning appears necessary to enable effective self-correction [mehandru2025bioagents, kamoi2024when].
Critique generation processes involve models analyzing own outputs and providing structured suggestions for improvement [madaan2023self]. Advanced critique frameworks implement multi-dimensional evaluation systems with carefully designed prompts instructing models to assess outputs based on specific criteria [ho2025self, madaan2023self, shinn2023reflexion]. Performance gains from self-refinement methods have been substantial including GPT-4 achieving gains of 8.7 points in code optimization and 13.9 points in code readability when using Self-Refine prompting [sahoo2024systematic].